# AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems

### v1

@@ Abstract
Feature selection is crucial in recommender systems for improving model efficiency and predictive performance. Traditional methods rely on agency models, such as decision trees or neural networks, to estimate feature importance. However, this approach is inherently limited, as the agency models may fail to learn effectively in all scenarios due to suboptimal training conditions (e.g., feature collinearity, high-dimensional sparsity, and data insufficiency). In this paper, we propose AltFS, an Agency-light Feature Selection method for deep recommender systems. AltFS integrates semantic reasoning from Large Language Models (LLMs) with task-specific learning from agency models. Initially, LLMs will generate a semantic ranking of feature importance, which is then refined by an agency model, combining world knowledge with task-specific insights. Extensive experiments on three public datasets from real-world recommender platforms demonstrate the effectiveness of AltFS. Our code is publicly available  [^1] The link will be released after publication for reproducibility.
AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems
Pengyue Jia1, Zhaocheng Du2, Yichao Wang2, Xiangyu Zhao1, Xiaopeng Li1, Yuhao Wang1, Qidong Liu1, Huifeng Guo2, Ruiming Tang2, 1City University of Hong Kong, 2Huawei Noah’s Ark Lab
1 Introduction
@@
Feature selection is crucial in deep recommender systems (DRS) for improving model performance, mitigating overfitting, and accelerating training and inference Wang et al. (2022); Lin et al. (2022). Numerous solutions have been proposed to address the need for feature selection in DRS, which can be broadly categorized into three types based on their selection techniques: (1) Shallow feature selection methods Breiman (2001); Chen and Guestrin (2016), which typically use statistical algorithms to assign importance scores to features. (2) Gate-based feature selection methods Guo et al. (2022); Lyu et al. (2023), where a gate vector is optimized during training by assigning gates to feature embeddings, and the gate values are interpreted as indicators of the corresponding features’ importance. (3) Sensitivity-based feature selection methods Wang et al. (2023); Zhang et al. (2023); Du et al. (2024), which evaluate parameter sensitivity using gradient information during backward propagation and then calculate feature importance.
Figure 1: Conventional paradigm and AltFS.
@@
As shown in Figure 1, all the aforementioned methods rely on training an agency model to capture the feature-to-label relationship. The accuracy of feature selection, therefore, hinges on how closely the agency model’s mapping aligns with the ground truth. However, in many real-world recommendation scenarios, agency models often fail to achieve optimal performance. For example, in cold-start or deep conversion tasks, the sparsity of samples—particularly positive samples—frequently leads to underfitting Schein et al. (2002). Conversely, in cases involving numerous high-cardinality features, agency models are prone to overfitting Zhang et al. (2022). Moreover, these models typically struggle to account for feature interdependencies, overlooking critical factors such as feature collinearity and complementarity when estimating feature importance.
@@
Figure 2: Overview of AltFS. There are three stages in AltFS, (1) Feature Importance Extraction Stage, (2) Feature Importance Refinement Stage, and (3) Retraining Stage.
LLMs offer potential solutions to the challenges faced by agency models. Trained on vast amounts of web data, LLMs can leverage their world knowledge to identify informative feature subsets in sparse scenarios and comprehend feature dependencies based on feature descriptions. For instance, agency models may fail to recognize that “longitude” and “latitude” must coexist to uniquely identify positions, potentially selecting only one for the feature subset. However, when this understanding is integrated into an LLM’s world knowledge, such issues can be easily resolved.
However, several challenges remain: (1) Feature Complexity: In real-world online platforms, numerous features exhibit complex relationships both among themselves and with the tasks at hand. (2) Knowledge Gap: LLMs are generally pre-trained on broad world knowledge, which may not align with the specific requirements of recommendation tasks. (3) Efficiency Demand: Additionally, industrial scenarios impose strict requirements on efficiency and resource consumption. Thus, efficiently and lightweightly integrating feature importance from world knowledge into recommendation model training poses a significant challenge.
@@
To address the challenges above, we propose AltFS, an Agency-light Feature Selection algorithm for DRS. Specifically, to explore complex feature interdependencies and feature-task relationships, we design a context-aware prompt iteration method that leverages LLMs to provide prior knowledge of feature importance. To efficiently and lightweightly integrate this knowledge into agency models for assessing feature importance, we introduce a bridge network to refine the feature importance rankings further. The bridge vector is trained under constraints to harmonize knowledge-based and task-specific feature importance. Our contributions can be summarized as follows:
@@
  • AltFS introduces the first paradigm that combines world knowledge priors with agency-based feature selection in DRS, effectively addressing traditional methods inaccuracies that arise from relying solely on the agency model.
@@
  • We design a prompt iteration method that enables LLMs to select effective features based on their world knowledge iteratively.
@@
  • We introduce a novel bridge network to integrate feature importance from world knowledge and the recommendation task space. This network is optimized in a lightweight, end-to-end manner.
@@
  • Extensive experiments are conducted on three public datasets derived from real-world online service platforms, verifying the effectiveness of our method.
2 Methodology
@@
In this section, we will detail AltFS. First, we provide an overview of the AltFS framework in Section 2.1. Next, we introduce the fundamental DRS model in Section 2.2 and the key modules of AltFS in Section 2.3 and Section 2.4. Finally, we will describe the optimization and retraining in Section 2.5 and Section 2.6.
2.1 Overview
@@
In this subsection, we provide an overview of AltFS, which consists of three stages: (1) Feature Importance Extraction, (2) Feature Importance Refinement, and (3) Retraining, as shown in Figure 2. In the Feature Importance Extraction stage, a prompt iteration method leverages LLMs to iteratively select predictive features from the candidate set, producing an initial feature importance ranking. To reduce bias, multiple LLMs are employed concurrently. Next, in the Feature Importance Refinement stage, the extracted feature importance sequences are refined using a bridge network integrated into the training process of the recommendation model. Finally, in the Retraining stage, the refined feature importance sequences are combined using the optimized bridge vector, generating a ranking that incorporates semantic reasoning and aligns with the recommendation task. The top $n$ features from this ranking are selected to retrain a recommendation model from scratch.
2.2 Basic DRS Architecture
2.2.1 Embedding Layer
@@
Efficient handling of categorical data is vital in recommender systems due to its inherent sparsity. Embedding layers address this challenge by transforming high-dimensional sparse categorical inputs into dense, low-dimensional vectors for DRS. The embedding process consists of two steps: (1) Binarization: Categorical features $\boldsymbol{x}$ are converted into binary vectors $\boldsymbol{x}^{\prime}$ using one-hot encoding, where the dimensionality depends on the number of unique values in each feature field. (2) Projection: The binary vectors are projected into dense embeddings via embedding table lookups. For each feature field, the binary vector $\boldsymbol{x}_{n}^{\prime}$ is mapped to an embedding $\boldsymbol{e}_{n}=\boldsymbol{x}_{n}^{\prime}\cdot\boldsymbol{M}_{n}$, where $\boldsymbol{M}_{n}$ is the learnable embedding table, and $\boldsymbol{e}_{n}\in\mathbb{R}^{dim}$ is the resulting embedding. The concatenated output is $\boldsymbol{E}=[\boldsymbol{e}_{1},\boldsymbol{e}_{2},\cdots,\boldsymbol{e}_{N}]$, where $\boldsymbol{E}$ represents the combined feature embeddings.
2.2.2 Transformation Layer
@@
$\boldsymbol{h}_{l}=\text{ReLU}(\boldsymbol{W}_{l-1}\boldsymbol{h}_{l-1}+\boldsymbol{b}_{l-1}),\ \boldsymbol{h}_{0}=\boldsymbol{E}$ (1)
$\hat{y}=\sigma(\boldsymbol{W}_{L}\boldsymbol{h}_{L}+\boldsymbol{b}_{L})$ (2)
@@
where $\boldsymbol{h}_{l}$ is the hidden state of the $l$th layer, and $\boldsymbol{W}_{l}$ and $\boldsymbol{b}_{l}$ are the weight and bias for the $l$th layer. For the output layer, $\boldsymbol{h}_{L}$, $\boldsymbol{W}_{L}$, and $\boldsymbol{b}_{L}$ denote the hidden state, weight, and bias, respectively. $\sigma$ represents the activation function for recommendation tasks (e.g., Sigmoid and Softmax).
2.3 Prompt Iteration
@@
To extract the complex semantic information between features and their relevance to specific recommendation tasks, we design the prompt illustrated in Figure 3, which consists of five parts: instructions, descriptions, feature sets, supplementary information, and output formatting. (1) Instructions: This section defines the role of LLMs and introduces the subsequent tasks. (2) Descriptions: This part provides essential information to address the feature selection problem, including descriptions of the recommendation task, dataset, and feature set. Specifically, in the recommendation task description, we outline the background of the task. For example, in the MovieLens dataset, the task is to determine whether a user likes a movie based on user profiles and movie information. The dataset description focuses on the number of samples and features. Finally, the feature set description is divided into user features, item features, and interaction features. Each feature description includes five elements: feature name, feature description, feature type (e.g., int, string), example, and unique values, providing sufficient information for LLMs to assess feature importance. (3) Feature Sets: The feature sets are composed of two parts: features already selected and candidate features. The union of these two sets constitutes the complete feature set. (4) Supplementary Information: This part emphasizes the expected actions of LLMs (selecting one feature from the candidate features to add to the already selected features during each inference) and provides reference criteria for feature selection. (5) Output Formatting: This section standardizes the output format of LLMs to facilitate automated and streamlined feature selection. The prompts we used in this work are detailed in Appendix A.1.
Figure 3: Structure of prompt.
Inspired by existing work Sun et al. (2023), we prompt LLMs to analyze and select informative features in an iterative manner. Considering that different LLMs may exhibit significant assessing bias due to variations in training data, we treat the LLM as an expert while simultaneously introducing multiple different LLMs to generate several feature importance sequences:
$f_{t}^{k}=\text{LLM}_{k}(\text{prompt},\boldsymbol{s}_{t}^{k},\boldsymbol{c}_{t}^{k})$ (3)
$\boldsymbol{s}_{t+1}^{k}=\boldsymbol{s}_{t}^{k}\cup\{f_{t}^{k}\},\ \boldsymbol{c}_{t+1}^{k}=\boldsymbol{c}_{t}^{k}\setminus\{f_{t}^{k}\}$ (4)
@@
where $\boldsymbol{s}_{t}^{k}$ and $\boldsymbol{c}_{t}^{k}$ denote the set of selected features and the set of candidate features at $t$th step with $k$th LLM, $f_{t}$ is the selected feature at $t$th step with $k$th LLM. The feature selected at $t$th step will be used to update the set of selected features and candidate features through Equation 4. Since user ID and item ID are essential in recommendation tasks, we initialize $\boldsymbol{s}_{0}$ as a set containing these two features. The iteration process continues until no features remain in the candidate feature set. The output is a matrix containing the features selected at each time step by different LLMs.
2.4 Bridge Network
The feature importance ranking obtained through prompting LLMs is based solely on general world knowledge, leading to a discrepancy with the feature importance derived during model training. To bridge this gap, we propose a bridge network that incorporates feature importance sequences produced by different LLMs into the training process of the recommendation model.
@@ 2.4 Bridge Network
For batch data, a few features will be first masked according to the feature importance ranking produced by LLMs. The number of masked features $r$ varies in each batch and is sampled from $r\sim\mathcal{U}(0,N\cdot\beta)$, where $N$ and $\beta$ are the number of feature fields and maximum mask ratio.
@@
where $\boldsymbol{\mathcal{M}}$ is the masking matrix with $m_{t}^{k}$ representing the masking value at $t$th step for $k$th LLM.
To integrate feature importance rankings produced LLMs into recommendation model training, we introduce a bridge vector. Suppose the feature set is $F$, $\boldsymbol{w}$ is the learnable bridge vector $\boldsymbol{w}=[w_{0},w_{1},\cdots,w_{K}]$ with a length equal to the number of LLMs, the integration process is:
$\delta(\boldsymbol{F}_{n},f_{t}^{k})=\begin{cases}1&\text{if}\ \boldsymbol{F}_{n}=f_{t}^{k}\\ 0&\text{otherwise}\end{cases}$ (6)
$\boldsymbol{\hat{w}}=\text{Softmax}(\boldsymbol{w}\cdot\text{exp}(\tau))$ (7)
$\boldsymbol{h}_{n}=\sum\nolimits_{k=0}^{K}\sum\nolimits_{t=0}^{N}\delta(\boldsymbol{F}_{n},f_{t}^{k})\cdot(m_{t}^{k}\cdot\boldsymbol{\hat{w}}_{k})$ (8)
where $\boldsymbol{F}_{n}$ is the $n$th feature field. $\boldsymbol{h}_{n}$ denotes the integrated weight for $n$th feature field. $K$ and $N$ denote the number of LLMs and feature fields, and $\tau$ is the temperature.
Before assigning the integrated weights $\boldsymbol{h}$ to embeddings, we need to normalize embeddings because their magnitude varies significantly across different feature fields, which can not be compared directly. We conduct BatchNorm to each field embedding by $\hat{\boldsymbol{e}}_{n}=(\boldsymbol{e}_{n}-\text{E}_{n})/\sqrt{\text{V}_{n}+\epsilon}$, where $\text{E}_{n}$ and $\text{V}_{n}$ are the mean and variance of $n$th feature field in a batch, $\epsilon$ is a constant introduces to enhance numerical stability. After transformation, the embeddings of features should be $\boldsymbol{\hat{E}}=[\boldsymbol{\hat{e}}_{1},\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{\hat{e}}_{N}]$. Next, we weights embeddings by $\boldsymbol{E}^{\prime}=[\boldsymbol{h}_{1}\boldsymbol{\hat{e}}_{1},\boldsymbol{h}_{2}\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{h}_{N}\boldsymbol{\hat{e}}_{N}]$, where $\boldsymbol{E}^{\prime}$ denotes weighted embeddings, $\boldsymbol{h}_{n}$ and $\boldsymbol{\hat{e}}_{n}$ are the weight and embeddings after transformation for $n$th feature field. The weighted embeddings will be input into the transformation layer in DRS to generate predictions following Equation 1 and Equation 2.
2.5 Optimization

### v2

# SELF: Surrogate-light Feature Selection with Large Language Models in Deep Recommender Systems

@@ Abstract.
Feature selection is crucial in recommender systems for improving model efficiency and predictive performance. Conventional approaches typically employ surrogate models—such as decision trees or neural networks—to estimate feature importance. However, their effectiveness is inherently constrained, as these models may struggle under suboptimal training conditions, including feature collinearity, high-dimensional sparsity, and insufficient data. In this paper, we propose SELF, an SurrogatE-Light Feature selection method for deep recommender systems. SELF integrates semantic reasoning from Large Language Models (LLMs) with task-specific learning from surrogate models. Specifically, LLMs first produce a semantically informed ranking of feature importance, which is subsequently refined by a surrogate model, effectively integrating general world knowledge with task-specific learning. Comprehensive experiments on three public datasets from real-world recommender platforms validate the effectiveness of SELF.
Feature Selection; Deep Recommender Systems; LLMs
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761378isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommender systems
1. Introduction
@@
Feature selection (Du et al., 2024b; Jia et al., 2024c) is essential in deep recommender systems (DRS) to enhance model performance (Wang et al., 2025), reduce overfitting, and accelerate both training and inference (Wang et al., 2022; Lin et al., 2022; Du et al., 2024a). A variety of approaches have been proposed to address this need, which can be broadly classified into three categories based on their selection strategies: (1) Shallow feature selection methods (Breiman, 2001; Chen and Guestrin, 2016), which typically use statistical algorithms to assign importance scores to features. (2) Gate-based feature selection methods (Guo et al., 2022; Lyu et al., 2023) optimize a gate vector during training by assigning learnable gates to feature embeddings. The resulting gate values are interpreted as indicators of feature importance. (3) Sensitivity-based feature selection methods (Wang et al., 2023; Zhang et al., 2023; Du et al., 2024a) assess parameter sensitivity using gradient information obtained during backpropagation, and subsequently compute feature importance.
Figure 1. Comparison of conventional paradigm and SELF.
@@
As illustrated in Figure 1, all the aforementioned methods rely on training a surrogate model to approximate the feature-to-label mapping. The effectiveness of feature selection critically depends on how well this surrogate model reflects the ground truth. However, in many real-world recommendation scenarios, surrogate models frequently fall short of delivering optimal performance. For instance, in cold-start or deep conversion tasks, the sparsity of samples—especially positive instances—often results in underfitting (Schein et al., 2002). In contrast, in scenarios characterized by numerous high-cardinality features, surrogate models are susceptible to overfitting (Zhang et al., 2022). Furthermore, these models commonly fail to capture interdependencies among features, thereby neglecting essential aspects such as collinearity and complementarity in the estimation of feature importance.
@@
Figure 2. Overview of SELF. There are three stages in SELF, (1) Feature Importance Extraction Stage, (2) Feature Importance Refinement Stage, and (3) Retraining Stage.
LLMs offer promising solutions to the aforementioned longstanding challenges. Trained on vast corpora of web data, LLMs possess extensive world knowledge that enables them to identify informative feature subsets even in sparse data settings and to capture complex feature dependencies using semantic feature descriptions. For example, traditional surrogate models may fail to infer that “longitude” and “latitude” must appear together to uniquely define a geographic location, often selecting only one as part of the feature subset. In contrast, LLMs, by virtue of their pretrained knowledge, can inherently understand such relationships and thereby mitigate these limitations.
Despite these advantages, several key challenges must still be addressed to fully leverage LLMs in this context: (1) Feature Complexity: In real-world online platforms, many features exhibit intricate interdependencies, both with each other and with the target task. (2) Knowledge Gap: While LLMs are pretrained on general-domain knowledge, this may not be directly aligned with the domain-specific requirements of recommendation systems or other downstream tasks. (3) Efficiency Demand: Moreover, industrial applications impose stringent constraints on computational efficiency and resource usage. Consequently, integrating feature importance derived from world knowledge into the training process of recommendation models in an efficient and lightweight manner remains a considerable challenge.
@@
To address the challenges above, we propose SELF, a SurrogatE-Light Feature selection method for DRS. Specifically, to explore complex feature interdependencies and feature-task relationships, we design a context-aware prompt iteration method that leverages LLMs to provide prior knowledge of feature importance. To efficiently and lightweightly integrate this knowledge into surrogate models for assessing feature importance, we introduce a bridge network to refine the feature importance rankings further. The bridge vector is trained under constraints to harmonize knowledge-based and task-specific feature importance. Our contributions can be summarized as follows:
@@
  • We propose SELF, the first paradigm that integrates world knowledge priors with surrogate-based feature selection in DRS, effectively mitigating the inaccuracies of traditional methods that rely solely on the surrogate model.
@@
  • We develop a prompt iteration strategy that enables LLMs to iteratively select effective features by leveraging their world knowledge.
@@
  • We present a novel bridge network that aligns feature importance derived from world knowledge with the recommendation task space. This network is optimized in a lightweight, end-to-end manner.
@@
  • We conduct extensive experiments on three public datasets sourced from real-world platforms, demonstrating the effectiveness of SELF.
2. Methodology
@@
In this section, we first provide an overview of the SELF framework in Section 2.1. Next, we introduce the fundamental DRS model in Section 2.2 and the key modules of SELF in Section 2.3 and Section 2.4. Finally, we will describe the optimization and retraining in Section 2.5 and Section 2.6.
2.1. Overview
@@
In this subsection, we present an overview of SELF, which comprises three stages: (1) Feature Importance Extraction, (2) Feature Importance Refinement, and (3) Retraining, as illustrated in Figure 2. In the Feature Importance Extraction stage, a prompt iteration strategy leverages LLMs to iteratively select predictive features from the candidate set, generating an initial feature importance ranking. To mitigate bias, multiple LLMs are employed in parallel. In the subsequent Feature Importance Refinement stage, the extracted importance sequences are refined via a bridge network integrated into the training process of the recommendation model. Finally, during the Retraining stage, the refined feature importance is fused through the optimized bridge vector, yielding a final ranking that incorporates semantic reasoning and aligns with the recommendation objective. The top-$d$ features from this ranking are then used to retrain the recommendation model from scratch.
2.2. Basic DRS Architecture
2.2.1. Embedding Layer
@@
Efficient handling of categorical data is vital in recommender systems due to its inherent sparsity. Embedding layers address this challenge by transforming high-dimensional sparse categorical inputs into dense, low-dimensional vectors for DRS. The embedding process consists of two steps: (1) Binarization: Categorical features $\boldsymbol{x}$ are converted into binary vectors $\boldsymbol{x}^{\prime}$ using one-hot encoding, where the dimensionality depends on the number of unique values in each feature field. (2) Projection: The binary vectors are projected into dense embeddings via embedding table lookups. For $n$-th feature field, the binary vector $\boldsymbol{x}_{n}^{\prime}$ is mapped to an embedding $\boldsymbol{e}_{n}=\boldsymbol{x}_{n}^{\prime}\cdot\boldsymbol{M}_{n}$, where $\boldsymbol{M}_{n}$ is the learnable embedding table, and $\boldsymbol{e}_{n}\in\mathbb{R}^{\text{dim}}$ is the resulting embedding. The concatenated output is $\boldsymbol{E}=[\boldsymbol{e}_{1},\boldsymbol{e}_{2},\cdots,\boldsymbol{e}_{N}]$, where $\boldsymbol{E}$ represents the combined feature embeddings.
2.2.2. Transformation Layer
@@
(1) $\boldsymbol{h}_{l}=\text{ReLU}(\boldsymbol{W}_{l-1}\boldsymbol{h}_{l-1}+\boldsymbol{b}_{l-1}),\ \boldsymbol{h}_{0}=\boldsymbol{E}$
(2) $\hat{y}=\sigma(\boldsymbol{W}_{L}\boldsymbol{h}_{L}+\boldsymbol{b}_{L})$
@@
where $\boldsymbol{h}_{l}$ is the hidden state of the $l$-th layer, and $\boldsymbol{W}_{l}$ and $\boldsymbol{b}_{l}$ are the weight and bias for the $l$-th layer. For the output layer, $\boldsymbol{h}_{L}$, $\boldsymbol{W}_{L}$, and $\boldsymbol{b}_{L}$ denote the hidden state, weight, and bias, respectively. $\sigma$ represents the activation function for recommendation tasks (e.g., Sigmoid and Softmax).
2.3. Prompt Iteration
To address the limitations of surrogate models in learning high-quality feature relationships, especially under conditions of feature sparsity or weak supervision, we introduce an external source of prior knowledge to guide the feature selection process. Specifically, we leverage the world knowledge embedded in LLMs to provide semantic priors that complement the surrogate model’s learning capacity. To this end, we design a prompt-based extraction strategy, as illustrated in the following box. The prompt comprises five components: instructions, descriptions, feature sets, supplementary information, and output formatting.
Prompt structure [Instructions] You are a professional researcher in recommender systems and feature selection. … [Descriptions] Task: <Task descriptions > Dataset: <Dataset descriptions > Features: <Feature descriptions > [Feature sets] Features already selected: <A set contains all selected features > Candidate features: <A set contains all candidate features > [Supplementary information] Please select one feature from the candidate feature set that you think is most important for this task to add to the selected features. Suppose we will discrete the features based on their unique values. The feature you choose should, as much as possible, have the following characteristics: 1. They are informative. 2. Independent of other selected features. 3. Simple and easy for the model to understand. [Output formatting] Your answer’s last line should be the feature name of your selected feature, without any other characters.
@@ 2.3. Prompt Iteration
(1) Instructions defines the role of LLMs and introduces the subsequent tasks. (2) Descriptions provide critical information for addressing the feature selection problem, encompassing the recommendation task, dataset, and feature set. For the task description, we outline the context and objectives of the recommendation scenario. For instance, in the MovieLens dataset, the task involves predicting whether a user will like a movie based on user profiles and movie information. The dataset description focuses on the number of samples and features. The feature set is categorized into user, item, and interaction features. Each feature is described using five attributes: name, description, data type (e.g., integer, string), example value, and the number of unique values. This structured representation offers sufficient context for LLMs to assess feature importance effectively. (3) Feature Sets are composed of two parts: features already selected and candidate features. The union of these two sets constitutes the complete feature set. (4) Supplementary Information emphasizes the expected actions of LLMs (selecting one feature from the candidate features to add to the already selected features during each inference) and provides reference criteria for feature selection. (5) Output Formatting standardizes the output format of LLMs to facilitate automated and streamlined feature selection. Inspired by existing work (Sun et al., 2023), we prompt LLMs to analyze and select informative features in an iterative manner. Considering that different LLMs may exhibit significant assessing bias due to variations in training data, we treat the LLM as an expert while simultaneously introducing multiple different LLMs to generate several feature importance sequences:
(3) $f_{t}^{k}=\text{LLM}_{k}(\text{prompt},\boldsymbol{s}_{t}^{k},\boldsymbol{c}_{t}^{k})$
(4) $\boldsymbol{s}_{t+1}^{k}=\boldsymbol{s}_{t}^{k}\cup\{f_{t}^{k}\},\ \boldsymbol{c}_{t+1}^{k}=\boldsymbol{c}_{t}^{k}\setminus\{f_{t}^{k}\}$
@@
where $\boldsymbol{s}_{t}^{k}$ and $\boldsymbol{c}_{t}^{k}$ denote the sets of selected and candidate features, respectively, at the $t$-th step for the $k$-th LLM, and $f_{t}^{k}$ represents the feature selected at that step. The selected feature $f_{t}^{k}$ is then used to update the selected and candidate sets according to Equation 4. Given that user ID and item ID are fundamental in recommendation tasks, we initialize $\boldsymbol{s}_{0}$ as a set containing these two features. The iterative process continues until the candidate feature set becomes empty. The final output is a matrix that records the features selected at each time step by different LLMs:
(5) $\boldsymbol{\mathcal{S}}=\begin{bmatrix}f_{0}^{0}&f_{1}^{0}&\cdots&f_{N}^{0}\\ f_{0}^{1}&f_{1}^{1}&\cdots&f_{N}^{1}\\ \vdots&\vdots&\ddots&\vdots\\ f_{0}^{K}&f_{1}^{K}&\cdots&f_{N}^{K}\\ \end{bmatrix}$
where each row corresponds to the feature selection trajectory of a specific LLM, and each column indicates the feature selected at a particular iteration step. Here, $K$ denotes the number of LLMs, and $N$ is the total number of selection steps.
2.4. Bridge Network
The feature importance rankings obtained via prompting LLMs rely exclusively on general world knowledge, which may diverge from the task-specific importance signals learned during model training. This discrepancy arises from a distributional mismatch between knowledge priors and empirical data. To effectively integrate knowledge-derived importance into the surrogate model’s learning process, we introduce a bridge network that acts as an intermediary, aligning the semantic priors from LLMs with the data-driven objectives of the recommendation model.
@@ 2.4. Bridge Network
For each training batch, we mask a subset of features according to the feature importance rankings produced by LLMs, beginning with the least important ones. The number of masked features, denoted by $r$, is sampled from a uniform distribution: $r\sim\mathcal{U}(0,N\cdot\beta)$, where $N$ is the total number of feature fields and $\beta$ is the maximum masking ratio.
@@
[…]
To incorporate the feature importance rankings from multiple LLMs into the training process, we introduce a learnable bridge vector that adaptively fuses their contributions. Let the feature set be denoted by $F$, and let $\boldsymbol{w}=[w_{0},w_{1},\dots,w_{K}]$ represent the bridge vector, where $K$ is the number of LLMs. The fusion weights are normalized via a temperature-scaled softmax:
(7) $\boldsymbol{\hat{w}}=\text{Softmax}(\boldsymbol{w}\cdot\exp(\tau))$
where $\tau$ is the temperature parameter controlling the sharpness of the softmax distribution.
For each feature field $\boldsymbol{F}_{n}$, we compute its final importance score $\boldsymbol{h}_{n}$ as follows:
(8) $\boldsymbol{h}_{n}=\sum_{k=0}^{K}\hat{w}_{k}\cdot\sum_{t=0}^{N}\mathbb{I}[f_{t}^{k}=\boldsymbol{F}_{n}]\cdot m_{t}^{k}$
where $\mathbb{I}[\cdot]$ is the indicator function that evaluates to 1 if feature $\boldsymbol{F}_{n}$ appears at step $t$ in the ranking of the $k$-th LLM, and 0 otherwise. $m_{t}^{k}$ is the masking value defined in Equation 6, indicating whether the feature at step $t$ in the $k$-th LLM’s sequence is retained. $\hat{w}_{k}$ denotes the normalized importance weight assigned to the $k$-th LLM. This formulation enables the model to dynamically weigh and aggregate unmasked feature importance across different LLM-generated rankings, producing a consolidated importance score $\boldsymbol{h}_{n}$ for each feature field.
Before applying the integrated weights $\boldsymbol{h}$ to the feature embeddings, we first normalize the embeddings to ensure consistency across feature fields. After transformation, the embeddings of features should be $\boldsymbol{\hat{E}}=[\boldsymbol{\hat{e}}_{1},\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{\hat{e}}_{N}]$. Next, we weight embeddings by $\boldsymbol{E}^{\prime}=[\boldsymbol{h}_{1}\boldsymbol{\hat{e}}_{1},\boldsymbol{h}_{2}\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{h}_{N}\boldsymbol{\hat{e}}_{N}]$, where $\boldsymbol{E}^{\prime}$ denotes weighted embeddings, $\boldsymbol{h}_{n}$ and $\boldsymbol{\hat{e}}_{n}$ are the weight and embeddings after normalization for $n$-th feature field. The weighted embeddings will be input into the transformation layer in DRS to generate predictions following Equation 1 and Equation 2.
2.5. Optimization